Consider a market with demand curve ?=200−? and suppose that the industry consists of a dominant firm which has a constant marginal cost equal to $40 per unit. There are ten other fringe producers; each has a marginal cost curve ??=40+10?,
where q is the output of a typical fringe producer. Assume there are no fixed costs for any producer.
a. What is the supply curve of the competitive fringe?
b. What is the dominant firm’s residual demand curve?
c. Find dominant firm’s profit-maximizing output and price. At this price, what is dominant firm’s market share?
Answer #1

MC = 40+100 P = 40 + 108. P- 40 - 108 8 = (-40) B = IP-40) X10 10 [8 = P-40) Fringe firm supply curve. (6) Dominant firm resi